OpenStudy (anonymous):

(24a^5+6a^3-9a^2+3a)/(3a)

OpenStudy (anonymous):

\[\large \frac{24a^5 +6a^3-9a^2+3a}{3a}=\frac{24a^5}{3a}+\frac{6a^3}{3a}-\frac{9a^2}{3a}+\frac{3a}{3a}\] Have fun cancelling out :)

OpenStudy (anonymous):